OpenStudy (anonymous):

A mass m is moving in a circular path of radius R. If the force causing the centripetal acc. of m, then what is the work done by F on m?

OpenStudy (anonymous):

\[Work = Force * Distance\] Which means that we only need distance for this equation to be complete. A great place to start is with arc length: \[s = \theta * r\]

OpenStudy (anonymous):

Here the Centripetal Force is always perpendicular to the displacement ,so no work is being done. Thus Work=0

OpenStudy (anonymous):

\[W = F*s*\cos \alpha if \alpha = 90 or \pi -> W = 0\]

OpenStudy (anonymous):

work done will be zero because the angle between acc and vel in circular is 90 degrees which makes work which is integration Force.ds equal to zero.The direction nof acc and force are same and same holds for vel and disp